Tech Staff Summary 12/1/00
Current Priorities in addition to routine tasks
Search for eastern hard rock site next week (also stop by PLAL).
Still need permissions on Edith and Library towers.
No word yet from Falcon.
Library will likely happen within next 45-60 days and we want to be ready to use their contracters to install and align.
Need to have at the very least MKTA, KOSE, and Richardson ready to align when Library is installed. Greg is working on subcontract for KOSE, LNXT, and NMAD.
Greg and Jim will get the final supplies needed.
MTTN hardware is installed and running
Walnut Mountain, Greentop, and Avondale Springs are complete.
Stays where it is. Need to patch holes in ceiling, add ventilation, yank non-functional shower, and install eye wash station.
This is becoming a high priority. We don't want to drag our feet but neither do we want to "burn" good field weather.
Will keep pluggin away as time permits.
Mitch still needs to complete the code to handle the gain channel. Subroutines are done and offline PII style is available. On the fly will need less robust than offline to avoid potentially large and time intensive buffering. (shall we start a pool on whether this ever gets done?)
Still need to get the battery charging room going. Who's had their brown recluse shots?
Have refurbished the USGS shack and Buddy has moved except for the junk in the fdl lab
Second tier priorities
Permit and install 18 ANSS strong motion sites by September 30, 2001
Andy has completed the painting. Jim will install a map mosaic covering areas where CERI has hardware
Mitch will produce a station map for the ANSS-MA region
Chris will install the white board
Greg will either fix or hide the nasty old map
Greg will learn that it is NOT a work bench
Cannabalize PI to make as much as PII operational.
Check with Arch or Jer-Ming for account numbers to buy supplies.
Not sure where the space exists to work on this.
EBZ/PWLA telemetry. Another repeater?
Secure propane tank.
replace coax jumpers
replace rg-58 unshielded jumpers
Need to record the SOH channel from the gps clocks to know when sync is lost (they will continue to produce IRIGE based on internal clock).
Greg has added channel to monitor gps clock. Need an ew module that monitors and complains.
Greg is finishing up planning for the I-40 bridge instrumentation.
On the horizon is planning for the AR borehole project.
Mitch and Steve will install earthworm v5 in June. This will include the
full blown schema capable of creating seed.
Status channels. Currently supplied on last 8 channels.
LNXT node permit/install L7 and L8
MKTA node permit/install remaining 3 ISIS